# Video-GroundingDINO: Towards Open-Vocabulary Spatio-Temporal Video Grounding

### v1

@@ Abstract
Video grounding aims to localize a spatio-temporal section in a video corresponding to an input text query. This paper addresses a critical limitation in current video grounding methodologies by introducing an Open-Vocabulary Spatio-Temporal Video Grounding task. Unlike prevalent closed-set approaches that struggle with open-vocabulary scenarios due to limited training data and predefined vocabularies, our model leverages pre-trained representations from foundational spatial grounding models. This empowers it to effectively bridge the semantic gap between natural language and diverse visual content, achieving strong performance in closed-set and open-vocabulary settings. Our contributions include a novel spatio-temporal video grounding model, surpassing state-of-the-art results in closed-set evaluations on multiple datasets and demonstrating superior performance in open-vocabulary scenarios. Notably, the proposed model outperforms state-of-the-art methods in closed-set settings on VidSTG (Declarative and Interrogative) and HC-STVG (V1 and V2) datasets. Furthermore, in open-vocabulary evaluations on HC-STVG V1 and YouCook-Interactions, our model surpasses the recent best-performing models by 4.26 m_vIoU and 1.83% accuracy, demonstrating its efficacy in handling diverse linguistic and visual concepts for improved video understanding. Our codes will be released at https://github.com/TalalWasim/Video-GroundingDINO.
@@ 4 Results
TubeDETR (CVPR’22) [28] VidSTG 29.88 46.73 19.37 51.63
STCAT (NeurIPS’22) [9] VidSTG 32.45 54.39 27.63 55.90
Video-GroundingDINO VidSTG 36.71 60.29 34.07 57.73
@@ 4.2.1 Open-Vocabulary Evaluation
Results on HC-STVG V1: We report open-vocabulary evaluation on m_vIoU, vIoU@0.3, and vIoU@0.5, along with the reduction in accuracy with respect to the closed-set supervised evaluation. We achieve state-of-the-art performance over both TubeDETR [28] and STCAT [9]. We attribute this strong performance to our design, which leverages the strong pre-trained generalized features of a spatial grounding foundation model.
@@ 4.2.2 Closed-Set Supervised Evaluation
Results on HC-STVG V1: We present results on the HC-STVG V1 dataset in Tab. 3, reporting m_vIoU, vIoU@0.3 and vIoU@0.5. We achieve a nearly $1.5$ unit gain in m_vIoU and vIoU@0.5 and a $1$ unit gain in vIoU@0.3 over the previous best method STVGFormer [11]. This shows the consistent performance of our method on this challenging dataset.
Results on HC-STVG V2: We present results on the HC-STVG V2 dataset in Tab. 4, reporting m_vIoU, vIoU@0.3 and vIoU@0.5. Our performance gain on HC-STVG V1 is reflected here as well, with a consistent performance gain on HC-STVG V2, in comparison to STVGFormer [11] and TubeDETR [28].
@@ 4.4 Limitation
While our video grounding model excels in closed-set and open-vocabulary scenarios, it leverages image-text pre-trained models like Grounding DINO [12]. To enhance understanding in open-vocabulary settings, an extension to video-language pre-training on a larger and more diverse dataset, akin to CLIP [15], can help further boost performance. Building a comprehensive video-language pre-training dataset with diverse natural language expressions and spatio-temporal localization is imperative, given the current constraints of datasets like VidSTG [32] and HC-STVG [24]. Future endeavors can explore creating a large-scale video-language grounding dataset, enabling models to learn generalized visual and textual representations for improved adaptation to out-of-distribution samples.
@@ 5 Conclusion
This paper introduces an Open-Vocabulary Spatio-Temporal Video Grounding task, addressing limitations in current methodologies by leveraging pre-trained spatial grounding models. The proposed model excels in closed-set and open-vocabulary scenarios, outperforming state-of-the-art results on datasets like VidSTG and HC-STVG. Featuring learnable adapter blocks for video-specific adaptation, it effectively bridges the semantic gap between natural language queries and visual content. This research contributes to understanding open-vocabulary challenges and lays a foundation for versatile video interpretation techniques.

### v2

# VideoGrounding-DINO: Towards Open-Vocabulary Spatio-Temporal Video Grounding

@@ Abstract
Video grounding aims to localize a spatio-temporal section in a video corresponding to an input text query. This paper addresses a critical limitation in current video grounding methodologies by introducing an Open-Vocabulary Spatio-Temporal Video Grounding task. Unlike prevalent closed-set approaches that struggle with open-vocabulary scenarios due to limited training data and predefined vocabularies, our model leverages pre-trained representations from foundational spatial grounding models. This empowers it to effectively bridge the semantic gap between natural language and diverse visual content, achieving strong performance in closed-set and open-vocabulary settings. Our contributions include a novel spatio-temporal video grounding model, surpassing state-of-the-art results in closed-set evaluations on multiple datasets and demonstrating superior performance in open-vocabulary scenarios. Notably, the proposed model outperforms state-of-the-art methods in closed-set settings on VidSTG (Declarative and Interrogative) and HC-STVG (V1 and V2) datasets. Furthermore, in open-vocabulary evaluations on HC-STVG V1 and YouCook-Interactions, our model surpasses the recent best-performing models by 4.88 m_vIoU and 1.83% accuracy, demonstrating its efficacy in handling diverse linguistic and visual concepts for improved video understanding. Our codes will be publicly released.
@@ 4 Results
TubeDETR (CVPR’22) [28] VidSTG 16.84 22.32 9.22 51.63
STCAT (NeurIPS’22) [9] VidSTG 22.58 32.14 20.83 55.90
VideoGrounding-DINO VidSTG 27.46 40.13 29.92 57.73
@@ 4.2.1 Open-Vocabulary Evaluation
Results on HC-STVG V1: We report open-vocabulary evaluation on m_vIoU, vIoU@0.3, and vIoU@0.5. We achieve state-of-the-art performance over both TubeDETR [28] and STCAT [9]. We attribute this strong performance to our design, which leverages the strong pre-trained generalized features of a spatial grounding foundation model.
@@ 4.2.2 Closed-Set Supervised Evaluation
Results on HC-STVG V1: We present results on the HC-STVG V1 dataset in Tab. 3, reporting m_vIoU, vIoU@0.3 and vIoU@0.5. We achieve a nearly $1.5$ unit gain in m_vIoU and vIoU@0.5 and a $1$ unit gain in vIoU@0.3 over the previous best method STVGFormer [11]. This shows the consistent performance of our method on this dataset.
Results on HC-STVG V2: We present results on the HC-STVG V2 dataset in Tab. 4, reporting m_vIoU, vIoU@0.3 and vIoU@0.5. Our performance gain on HC-STVG V1 is reflected here as well, with a consistent performance gain on HC-STVG V2, in comparison to the SoTA [11, 28].
@@ 4.4 Limitation
While our video grounding model excels in closed-set and open-vocabulary scenarios, it leverages image-text pre-trained models like Grounding DINO [12]. To enhance understanding in open-vocabulary settings, an extension to video-language pre-training on a larger and more diverse dataset, akin to CLIP [15], can help further boost performance. Building a video-language pre-training dataset with diverse natural language expressions and spatio-temporal localization is imperative, given the constraints of datasets like VidSTG [32] and HC-STVG [24].
@@ 5 Conclusion
This paper introduces an Open-Vocabulary Spatio-Temporal Video Grounding task, enhancing current closed-set methodologies by using pre-trained representations from spatial grounding models. The proposed model performs well in closed-set and open-vocabulary scenarios, surpassing state-of-the-art results in supervised setting on VidSTG and HC-STVG datasets, and outperforming recent models in open-vocabulary on HC-STVG V1 and YouCook-Interactions. Its architecture includes learnable adapter blocks for video-specific adaptation, bridging the semantic gap between natural language queries and visual content. This research addresses open-vocabulary challenges and explores achieving robust performance without extensive video annotations, paving the way for open-vocabulary video grounding.